OpenStudy (anonymous):

Help me pleaseeeee An initial population of 400 deer decreases at an annual rate of 9%. Write an exponential function to model the deer population. What will the approximate population be after 3 years? f(x) = (400 times 0.09)^x; 46,656 f(x) = 400 (0.91)^x; 301 f(x) = 400 (1.09)^x; 518 f(x) = 400 (9)^x; 301

OpenStudy (anwaarullah):

400(1.09)^x

OpenStudy (anwaarullah):

An exponential growth or decay function is a function that grows or shrinks at a constant percent growth rate. The equation can be written in the form x f (x) = a(1+ r) or x f (x) = ab where b = 1+r Where a is the initial or starting value of the function r is the percent growth or decay rate, written as a decimal b is the growth factor or growth multiplier. Since powers of negative numbers behave strangely, we limit b to positive values.

OpenStudy (anonymous):

Omg thank youuuuuuuuu

OpenStudy (anwaarullah):

welcom lol